PLease please help me with z-scores!
Set up:
The speed of automobiles on a highway during rush hour is normally distributed with a mean of 62mph and standard deviation of 5mph.
The question:
Label the normal distribution with the mean and the values corresponding to z= 1, 2, 3, -1, -2, -3
I found the scores for the values but I feel like I did it wrong because I didn't do any calculation with the mean and I have no clue where to start. please help

Answer by rothauserc(4718)

if z = 1, then
:
z-score = (X - mean)/standard deviation
:
1 = (X - 62)/5
:
X - 62 = 5
:
X = 67
:
if z = -1
:
-1 = (X -62)/5
:
X - 62 = -5
:
X = 57
:
the rest of the X values follow what I did above
: